OpenStudy (anonymous):

The legs of a right triangle are 18 cm and 80 cm. What is the length of the hypotenuse? 82 cm 85 cm 89 cm 98 cm

OpenStudy (ali2x2):

this is simple .-.

OpenStudy (anonymous):

@ganeshie8 @pooja195 @nincompoop @heretohelpalways @ali2x2

OpenStudy (anonymous):

98?

OpenStudy (ali2x2):

god @heretohelpalways STAPH STEALING EMH COOL I WAS HELPING HIM >:o

OpenStudy (ali2x2):

a^2 + b^2 = c^2 so 18^2 + 80^2 = c^2 324 + 6400 = c^2 6724 = c^2 c = square root of 6724

OpenStudy (anonymous):

lol, okay I will stop

OpenStudy (ali2x2):

lol ok xD

OpenStudy (anonymous):

is it 98?

OpenStudy (anonymous):

@ali2x2 @heretohelpalways

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

find the square root of 6724

OpenStudy (ali2x2):

82 cm a(2) + b(2) = c(2) 18(2) + 80(2) = c(2) 324 + 6400 = c(2) 6724 = c(2)

OpenStudy (ali2x2):

then you get what?

OpenStudy (ali2x2):

82

OpenStudy (anonymous):

thank you

OpenStudy (ali2x2):

mhm